The theory of orthogonal polynomials and spectral analysis of Schrödinger operators.

Serguei Denissov California Institute of Technology

Some new results in the spectral analysis of Schrödinger operators will be presented. Deep relations to the theory of polynomials orthogonal on the unit circle will be discussed.